how many grams of water would be produced when 1.0 moles of h2 gas reacts with excess o2 gas


1
Expert's answer
2021-03-22T06:36:29-0400

H2 (g)+ 1/2 O2 (g)"\to" H2O

1 mol of H2 gas react with excess O2 gas

Mass of 1 mole of H2 gas = 2.01565 g

Mass of 1/2 mole of O2 gas = (1/2) × 32g = 16g

Water produced = 2.01565 g + 16 g

= 18.01565 g